OpenStudy (anonymous):

help solve the sum of three consecutive integers is -123

OpenStudy (anonymous):

40 41 42

OpenStudy (anonymous):

negatives

OpenStudy (anonymous):

Let the three consecutive integers be x, (x + 1) and (x + 2) So, the question says their sum is -123 So, x + x +1 x + 2 = -123 3x = -123 - 3 x = -126 /3 = -42 So the numbers are -42, -41 and -40...